Ama is now 6 times as old as kofi. In 12years time ama will be 2 times as old as kofi. How old are they.
Answer by Alan3354(69443)

Here's a similar problem:
======================
5 years ago I was 5 times my son's age, in 8 years time I will be 3 times his age. How old am I today?
==============
F = Father's age, S = Son's age
----
F - 5 = 5*(S - 5) = 5S - 15
F + 8 = 3*(S + 8) = 3S + 24
===============
F - 5 = 5S - 15
F + 8 = 3S + 24
---
Solve for F & S